# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ A.3. Acronyms
AI-PAML Adaptive Incremental Passive-Aggressive Machine Learning
AIS Automatic Identification System
@@
ANRS Advanced Navigation Research Simulator
@@
BIMCO Baltic and International Maritime Council
DAU Data Acquisition Unit
DDoS Distributed Denial of Service
DoS Denial of Service
DRFM Digital Radio Frequency Memory
DSIAS Digital Signature Based Identity Authentication Scheme
EA Electronic Attack
ECDIS Electronic Chart Display and Information System
ECDSA Elliptic Curve Digital Signature Algorithm
eMIR eMaritime Integrated Reference Platform
ENC Electronic Navigational Chart
FMEA Failure Mode and Effect Analysis
FMECA Failure Mode, Effect, and Criticality Analysis
GHG Greenhouse Gas
GP Gaussian Process
GIS Geographic Information System
GMDSS Global Maritime Distress and Safety System
GNSS Global Navigation Satellite System
GPS Global Positioning System
HAZOP Hazard and Operability Analysis
HFACS Human Factors Analysis and Classification System
HW-F High Weight Function
IACS International Association of Classification Society
IBS Integrated Bridge System
IMO International Maritime Organisation
IoT Internet of Things
ISPS International Ship and Port Facility Security
IT Information Technology
lite-CA Lite Certification Authority
LSTM Long Short-Term Memory
Acronym Description
@@
MARL Multi-Agent Reinforcement Learning
@@
MCAD Maritime Cyber Attack Database
mIBC Maritime Certificate-less Identity-Based Public Key Cryptography
@@
MTPPS Mix-zone Based Trajectory Privacy Protection Scheme
NAVTEX Navigation Telex
@@
NIST National Institute of Standards and Technology
NMEA National Marine Electronics Association
OT Operational Technology
PDCA Plan-Do-Check-Act
PKI Public Key Infrastructure
PLA Port of London Authority
@@
PLCs Programmable Logic Controllers
PoC Proof-of-Concept
PPI Plan Position Indicator
RSSI Received Signal Strength Indicator
S2ML Dual Stack Machine Learning
SATCOM Satellite Communications
SEI Specific Emitter Identification
RADAR RAdio Detection And Ranging
UAV Unmanned Aerial Vehicle
@@
VDES VHF Data Exchange System
VDR Voyage Data Recorder
VHF Very High Frequency
@@
VSAT Very Small Aperture Terminal

### v3

@@ A-C Acronyms
AI-PAML Adaptive Incremental Passive-Aggressive Machine Learning ISPS International Ship and Port Facility Security
AIS Automatic Identification System IT Information Technology
@@
ANRS Advanced Navigation Research Simulator lite-CA Lite Certification Authority
@@
BIMCO Baltic and International Maritime Council LSTM Long Short-Term Memory
@@
DAU Data Acquisition Unit MARL Multi-Agent Reinforcement Learning
@@
DDoS Distributed Denial of Service MCAD Maritime Cyber Attack Database
@@
DoS Denial of Service MTPPS Mix-zone Based Trajectory Privacy Protection Scheme
DRFM Digital Radio Frequency Memory NAVTEX Navigation Telex
@@
DSIAS Digital Signature Based Identity Authentication Scheme NIST National Institute of Standards and Technology
EA Electronic Attack NMEA National Marine Electronics Association
ECDIS Electronic Chart Display and Information System OT Operational Technology
ECDSA Elliptic Curve Digital Signature Algorithm PDCA Plan-Do-Check-Act
eMIR eMaritime Integrated Reference Platform PKI Public Key Infrastructure
ENC Electronic Navigational Chart PLA Port of London Authority
@@
FMEA Failure Mode and Effect Analysis PLCs Programmable Logic Controllers
FMECA Failure Mode, Effect, and Criticality Analysis PoC Proof-of-Concept
GHG Greenhouse Gas PPI Plan Position Indicator
GP Gaussian Process RSSI Received Signal Strength Indicator
GIS Geographic Information System S2ML Dual Stack Machine Learning
GMDSS Global Maritime Distress and Safety System SATCOM Satellite Communications
GNSS Global Navigation Satellite System SEI Specific Emitter Identification
GPS Global Positioning System RADAR RAdio Detection And Ranging
HAZOP Hazard and Operability Analysis UAV Unmanned Aerial Vehicle
@@
HFACS Human Factors Analysis and Classification System VDES VHF Data Exchange System
HW-F High Weight Function VDR Voyage Data Recorder
IACS International Association of Classification Society VHF Very High Frequency
@@
IBS Integrated Bridge System VSAT Very Small Aperture Terminal
Appendix B Biography Section
[Uncaptioned image] Meixuan Li received the B.E. degree (Hons.) in computer science from Singapore University of Technology and Design (SUTD), Singapore, in 2022. After graduation, she works as a Research Officer for iTrust at SUTD. Her work is primarily focused on the Maritime Testbed of Shipboard Operational Technology (MariOT) project, a strategic initiative funded through the National Satellite of Excellence (NSoE) program by the Singapore government. Her research interests include cyber-physcial security, operational technology system, maritime cybersecurity, and autonomous shipping.
[Uncaptioned image] Jianying Zhou is a professor and center director for iTrust at Singapore University of Technology and Design (SUTD). He received PhD in Information Security from Royal Holloway, University of London. His research interests are in applied cryptography and network security, cyber-physical system security, mobile and wireless security. He has published papers at international conferences (IEEE S&P, ACM CCS, USENIX Security, NDSS etc.) and journals (IEEE TIFS, IEEE TDSC etc.). He received ESORICS’15 Best Paper Award and ACSAC’23 Distinguished Paper Award. He is a co-founder & steering committee co-chair of ACNS, steering committee chair of ACM AsiaCCS, and steering committee member of Asiacrypt. He is associate editor-in-chief of IEEE Security & Privacy. He is also an ACM Distinguished Member. He received the ESORICS Outstanding Contribution Award in 2020, in recognition of his contributions to the community.
[Uncaptioned image] Sudipta Chattopadhyay received the PhD degree in computer science from the National University of Singapore, Singapore, in 2013. He is an associate professor with the Information Systems Technology and Design Pillar, Singapore University of Technology and Design, Singapore. In his doctoral dissertation, he researched on Execution-Time Predictability, focusing on Multicore Platforms. He seeks to understand the influence of execution platform on critical software properties, such as performance, energy, robustness, and security. His research interests include program analysis, embedded systems, and compilers. His serves in the review board of the IEEE Transactions on Software Engineering.
[Uncaptioned image] Mark Goh has a B.Eng in Civil Engineering (Hons.) and an M.Eng in Environmental Science and Engineering (Hons.) from the National University of Singapore (NUS). He started his career at the Maritime and Port Authority of Singapore (MPA), managing R&D projects and programmes in shipboard technology, renewable energy and civil and structural engineering. Mark moved on to develop stakeholder engagement and partnership strategies for a diversity of stakeholders at the national level. At iTrust, he oversees international collaborations, office and lab administration, and business development, securing research funds and revenue for the Centre.